OpenStudy (anonymous):

In right triangle ABC, ∠A and ∠B are complementary angles and sin A = 89. What is cos B? 817−−√17 89 17−−√9 17−−√8

OpenStudy (anonymous):

this doesnt pertain to music

OpenStudy (chantysquirrel1129**):

Why is your name "239iabusekids"?

OpenStudy (anonymous):

Complementary angles are two angles that add up to a measure of 90 So, m<a+m<b=90 89+m<b=90 90-89=m<b 1=m<b So the measure of angle B is 1 Therefore none of those answers are correct.